Grade 12Electric Current

State how the null condition of the bridge will change when the position of the battery and the galvanometer used in a Wheatstone bridge circuit are altered mutually.

raja
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
At the point when the position of the battery and the galvanometer utilized as a part of a Wheatstone connect circuit are changed commonly. As the wheatstone connect states that 
P/Q = R/S = l/100-l 
After trade the position of the battery and the galvanometer are changed commonly. 
P/Q = R/S = 100-l/l